OpenStudy (francis1):

Convert3,284 to appropriate metric units

OpenStudy (mathstudent55):

3,284 what? The unit is missing?

OpenStudy (francis1):

it doesnt say

OpenStudy (anonymous):

is it miles

OpenStudy (francis1):

yes

OpenStudy (anonymous):

You most likely will want to convert to kilometers. 1 mile = 1.61 kilometers

OpenStudy (anonymous):

5285.09 km

OpenStudy (francis1):

huh?

OpenStudy (anonymous):

can you help me???? pleanse and thank you

OpenStudy (anonymous):

When you convert 3,284 Miles . It becomes a 5285.09 km.

OpenStudy (mathstudent55):

1 mile = 1.609 km Divide both sides by 1.609 km to get (1 mile)/(1.609 km) = 1 If you divide the top equation by 1 mile, you get 1 = (1.609 km)/(1 mile) Since you have two ratios equal to 1, they are equal: (1 mile)/(1.609 km) = (1.609 km)/(1 mile)

OpenStudy (mathstudent55):

Now look at your original problem. You need to convert miles to km. 3,284 miles, when multiplied by (1.609 km)/(1 mile), the mile units will cancel leaving km. So the conversion is: 3,284 miles * (1.609 km)/(1 mile) = Just do the math

OpenStudy (anonymous):

@Jtgovernor its tan33=x/44 then it will become 44tan33=x......x=28.6 ft

OpenStudy (anonymous):

thank you!!!!!!! :)

OpenStudy (mathstudent55):

For the angle of 44 deg, the side labeled x is the opposite leg. The side labeled 44 is the adjacent leg. The trig function that relates the opposite to the adjacent is the tanglent. tanA = opp/adj, so you have tan 33 = x/44 Solve for x

OpenStudy (francis1):

what about 21 inches

OpenStudy (anonymous):

on more????? and thank you so much!!!!!

OpenStudy (mathstudent55):

@Jtgovernor Why are you posting new problems in someone else's questions? This should be one post for each question. Start a new post and I'll help you there.

OpenStudy (anonymous):

@mathstudent55 I already helped him.

OpenStudy (anonymous):

thanks @Shames

OpenStudy (mathstudent55):

@francis1 21 inches converted to metric. The metric unit that is comparable with the inch is the cm. The conversion factor is 2.54 cm = 1 in.

OpenStudy (anonymous):

@francis1 53.34 cm

OpenStudy (mathstudent55):

With this you get two fractions, both equal to 1: (1 in.)/(2.54 cm) = (2.54 cm)/(1 in.) = 1

OpenStudy (mathstudent55):

Since you have 21 inches, you need to cancel inches, so you need the fraction that has inches in the denominator. 21 in. * (2.54 cm)/( 1 in.) = 53.34 cm

OpenStudy (francis1):

how about 1 pint

OpenStudy (anonymous):

yes please

OpenStudy (mathstudent55):

A pint is comparable to a liter. In fact, a quart is very close to a liter, but a liter can also be used for pints. 1 pint = 0.473176 liter

OpenStudy (mathstudent55):

Once again, we have (1 pint)/(0.473176 liter) = (0.473176 liter)/(1 pint) = 1 Since you have a pint, and want to get rid of it, multiply by the conversion factor that has pints in the denominator. 1 pint * (0.473176 liter)/(1 pint) = 0.473176 liter

OpenStudy (francis1):

325 f

OpenStudy (mathstudent55):

This is a different story. To convert from deg C to deg F or vice versa, you need a formula. C = (5/9)(F - 32) Plug in the degrees F into F in the formula, and the degrees C with be C in the formula.

OpenStudy (francis1):

ok thank you very much for all your help hope some day i can helpu this much and more